how do u solve a system by substitution x+y=14 and -7x+7y=-14

Answer by Theo(13342)

x + y = 14
-7x + 7y = -14

Solve for y in the first equation to get:

y = 14 - x

Substitute for y in the second equation to get:

-7x + 7 * (14-x) = -14

Solve for x.

Remove Parentheses to get:

-7x + 7*14 - 7x = -14

Combine like terms to get:

-14x + 7*14 = -14

subtract 7*14 from vboth sides to get:

-14x = -14 - 7*14

Simplify and combine like terms to get

-14x = -112

Divide both sides by -14 to get:

x = 8
Now that you have x = 8, go back to your first equation and solve for y

Your first equation is:

x + y = 14

Replace x with 8 to get:

8 + y = 14

Solve for y to get y = 6

You have:

x = 8
y = 6

Substitute in second equation to get:

-7*8 + 7*6 = -14

This becomes -56 + 42 = -14 which becomes -14 = -14 which is true.

Values of x = 8 and y = 6 are good.